OpenStudy (anonymous):

What is the value of the coefficient of "a" when the quadratic equation y = (3x + 1)(3x − 2) is written in standard form?

OpenStudy (anonymous):

if the equation is written as f(x)=ax^2+bx+c then a = 9 but if the equation is written as f(x)=cx^2+bx+a then a = -2

OpenStudy (anonymous):

this equation in the standard form is f(x)= 9x^2 - 3x -2

OpenStudy (anonymous):

okayy thankyouu